Solve the following recurrence relation
a) an = 3an-1. n≥1
b) an = 5an-1 + 2 n≥1
c) an = 7an-1 + n n≥1
1
Expert's answer
2020-11-17T16:56:30-0500

a) "a_n = 3a_{n-1}, n\u22651"


The characteristic equation is "\lambda =3", and therefore the solution is "a_n=C\cdot 3^n" where "C=a_0."


b) "a_n = 5a_{n-1} + 2, n\u22651"


The characteristic equation of "a_n = 5a_{n-1}" is "\lambda =5", and therefore the solution is "a_n=C\cdot 5^n+b_n" where "b_n = b" is a solution of "a_n = 5a_{n-1} + 2". Therefore, "b=5b+2" which is equivalent to "b=-\frac{1}{2}" . So, "a_n=C\cdot 5^n-\frac{1}{2}." If "n=0" then "a_0=C-\frac{1}{2}", and thus "C=a_0+\frac{1}{2}." Consequently, the solution is "a_n=(a_0+\frac{1}{2})\cdot 5^n-\frac{1}{2}".


c) "a_n = 7a_{n-1} + n, n\u22651"


The characteristic equation of "a_n = 7a_{n-1}" is "\lambda =7", and therefore the solution is "a_n=C\cdot 7^n+b_n" where "b_n = an+b" is a solution of "a_n = 7a_{n-1} + n". Therefore, "an+b=7(a(n-1)+b)+n" which is equivalent to "(6a+1)n+6b-7a=0". So, "6a+1=0" and "6b-7a=0" which is equivalent to "a=-\frac{1}{6}" and "b=-\frac{7}{36}." Therefore, "a_n=C\cdot 7^n-\frac{1}{6}n-\frac{7}{36}." If "n=0" then "a_0=C-\frac{7}{36}", and thus "C=a_0+\frac{7}{36}." Consequently, the solution is "a_n=(a_0+\frac{7}{36})\cdot 7^n-\frac{1}{6}n-\frac{7}{36}."